# Boltzmann-Shannon Index: A Geometric-Aware Measure of Clustering Balance

### v2

@@ Abstract
We introduce the Boltzmann–Shannon Index (BSI), a normalized measure for clustered continuous data that captures the interaction between frequency-based and geometry-based probability distributions. Building on ideas from geometric coarse-graining and information theory, the BSI quantifies how well a partition reflects both the population of each cluster and its effective geometric extent. We illustrate its behavior on synthetic Gaussian mixtures, the Iris benchmark, and a high-imbalance resource-allocation scenario, showing that the index provides a coherent assessment even when traditional metrics give incomplete or misleading signals. Moreover, in resource-allocation settings, we demonstrate that BSI not only detects severe density–geometry inconsistency with high sensitivity, but also offers a smooth, optimization-ready objective that naturally favors allocations balancing demographic weight with each group’s effective spread in the outcome space, while providing a smooth, gradient-friendly regularizer that can be easily embedded in modern policy-making and algorithmic governance optimization frameworks.
@@ I Introduction
Information theory is the mathematical study of the quantification, storage and communication of information. Largely credited with adapting the concept of entropy to a new context, Claude Shannon is widely regarded as the father of information theory. In 1948, he established a framework for understanding the fundamental limits of signal processing and data compression, which included quantifying channel capacity shannon1948mathematical. However, as stated by Max Planck in 1900, “The logarithmic connection between entropy and probability was first stated by Ludwig Boltzmann in his kinetic theory of gases.” planck1914theory
In fact, in his pioneering work, Boltzmann laid the foundation for statistical mechanics by introducing entropy as a measure of disorder arising from the microscopic configurations of a physical system boltzmann1866mechanische. It was almost a century later that Shannon reframed entropy in the context of information theory, defining it as a quantitative measure of uncertainty in symbolic communication.
While Shannon’s formulation has become the cornerstone of modern data science, it is intended for discrete symbolic states and all applications to continuous states involve assuming probability mass functions shannon1948mathematical. As Jaynes emphasized in his foundational reformulation of information theory through the lens of statistical mechanics jaynes1957information, this assumption is not "innocuous": Shannon entropy is always defined relative to an underlying measure, and treating discrete categories as equally likely corresponds to choosing a uniform prior over the state space:
@@
When analyzing data in continuous state spaces, frequency of observation is not the only thing that matters; another important factor to consider is the geometry of the data. Data points occupy space, form shapes, and concentrate in regions of varying density. For this reason, and in an effort to address this gap, recent work focused on a form of metric entropy sinai1959notion; kolmogorov1985new, termed geometric partition entropy diggans2022geometric; diggans2025generalizing, that defines entropy directly from the spatial organization of data rather than from symbolic frequencies. Geometric partition entropy (GPE), instead of using traditional histogram binning, partitions the data into coarse-grained regions so that each bin contains approximately the same number of sample points (e.g., using quantiles in one dimension), providing a geometry-driven measure of uncertainty.
@@
Figure 1: Comparison between frequency-based and geometry-based quantifications of uncertainty in one-dimensional data. A representative picture of the underlying probability distribution $\rho(x)$ illustrates how data occupy the state space (a), together with the corresponding cumulative distribution function $CDF(x)$ (b). A traditional, frequency-based estimate approximates the distribution by evaluating differences in the empirical CDF at fixed bin boundaries (c); the probability $p_{i}$ of the $i$-th bin is given by the height of the corresponding blue step. In contrast, geometric partition entropy (GPE) constructs a measure-based representation by partitioning the probabilities on $[0,1]$ into equal masses and mapping these partitions back through the inverse CDF (d); the resulting geometric lengths $l_{i}$ define the bin sizes as proportions of the total state-space length. Figure adapted with permission from the authors of diggans2023boltzmann.
@@
For many complex data distributions, including the mixed continuous/discrete cases that arise in machine learning applications, there is clear utility in both views, and so further development is needed for information–theoretic approaches that respect both where data lie in space (geometry-based probability, $q$) and how often they occur in particular regions (frequency-based probability, $p$). Building on previous notions of the importance of geometric distribution to entropy kozachenko1987sample, the Boltzmann-Shannon Interaction Entropy (BSIE) diggans2023boltzmann continued in that vein. By combining the Boltzmann-inspired geometric view of GPE with Shannon’s frequentist formulation, BSIE provided a normalized measure that captures how the underlying geometry of continuous variables interacts with the estimated probability mass function; rather than treating frequency and shape as separate aspects of a dataset, BSIE quantifies their joint structure.
@@
But, while BSIE provided a measure of the interaction between frequency and geometry in continuous variables, it did not directly account for scenarios in which the data are already partitioned (i.e., clustered). In these common situations, the challenge is not necessarily to quantify the entropy of a distribution, but to understand the impact that a given choice of partition has on the resulting symbolic representation of the dynamics through the underlying geometry and density of the states.
@@
For example, it is common practice to label a dataset as imbalanced solely on the basis of the target distribution, typically when one class accounts for the vast majority of observations. However, especially when data lie in continuous spaces, we will show that imbalance is not determined by frequency-based probability alone. Consider a scenario where the outcome variable follows a $90:10$ distribution, but the underlying state space is also partitioned in a way that reflects the same $90:10$ proportion. In such a case, the dataset should not be considered imbalanced; the frequency distribution is fully consistent with the geometric structure of the data. This illustrates that true balance or imbalance arises from the interaction between class frequencies and the geometry of the feature space, not from frequencies in isolation.
Here, we introduce a measure of how well a dataset is spread across clusters that define coarse states having both a frequency-based probability and a geometric-based probability; in other words, we will introduce an index, which we will refer to as the Boltzmann-Shannon Index (BSI), that will give a sense of how well the state-space is partitioned, with equal densities among the clusters resulting in the ideal score of unity.
@@
Unlike existing clustering validity indices, the proposed Boltzmann–Shannon Index captures both the geometric spread of the data and the frequency with which regions of the state space are occupied. Distance-based metrics, such as the Silhouette Score rousseeuw1987silhouettes, Davies–Bouldin Index davies2009cluster, and Calinski–Harabasz Index calinski1974dendrite, quantify compactness or separation but rely solely on pairwise distances or variance ratios, ignoring how probability mass is distributed across clusters; moreover, they are not normalized and can be difficult to compare across datasets or scales. Information-theoretic measures, such as Normalized Mutual Information (NMI) strehl2002cluster, overcome some of these issues but require ground-truth labels and therefore cannot be used in fully unsupervised settings. In contrast, the Boltzmann–Shannon Index is agnostic, normalized, and uniquely incorporates both geometry-based and frequency-based probabilities, enabling it to detect imbalances in how the state space is partitioned that traditional metrics cannot reveal.
@@
The remainder of the paper is structured as follows. In Section II we review the basics and importance of data clustering. In Section III we define the Boltzmann-Shannon Index, while in IV we apply this measure to several data sets and comment on interesting use-cases and future work. We conclude with a summary of the work and plans for adaptation to mixed continuous/discrete datasets.
II Clustering
Since our index is designed to evaluate datasets that have already been clustered or otherwise partitioned, we begin with a brief overview of common clustering approaches. For the purposes of this work, the specific algorithm used is irrelevant as the BS-Index operates "post-hoc" on any set of cluster labels. Its role is to quantify how effectively a given clustering or partitioning scheme organizes a continuous state space into a meaningful coarse-grained representation.
Since our interest lies in evaluating a given partition rather than in how it is produced, we limit our overview to only the main clustering methods, some of them used as motivating examples. These include the $K$-Means lloyd1982least, which partitions data into $K$ clusters by minimizing within-cluster variance, Spectral Clustering shi2000normalized; ng2001spectral, which leverages the eigenstructure of a similarity graph to identify nonlinearly separable groups, and DBSCAN (Density-Based Spatial Clustering of Applications with Noise) ester1996density, a density-based method that discovers clusters as connected regions of high point density. We do not provide formal definitions here and instead refer the interested reader to standard references for full methodological details.
III Methodology
@@
Given a data set $\mathcal{X}=\left\{x_{0},x_{1},...,x_{N}\right\}\subset\mathbb{R}^{d}$ and labels $\mathcal{L}=\left\{\ell_{0},\ell_{1},...,\ell_{N}\right\}$ where $\ell_{i}\in\left\{1,...,K\right\}$, we define the Boltzmann-Shannon Index to be one minus the Jensen Shannon Divergence between a geometric measure-based distribution over the $K$ states of the partition, i.e. cluster, and the frequency-based distribution of the same $K$ states. If we let $q$ represent a normalized geometric distribution and $p$ represent the more commonly used normalized frequency-based distribution, then we define the BS-index as:
$BS_{Index}$ $=1-JSD(p||q)$ (1)
$=1-\frac{D_{KL}(p||m)+D_{KL}(q||m)}{2},$
@@
where $D_{KL}(x|y)$ is the Kullback-Leibler divergence between distributions $x$ and $y$, and $m$ is taken to be the midpoint distribution defined as the average of $p$ and $q$, i.e. $m=(p+q)/2$.
@@
Although the Kullback-Leibler divergence is asymmetric, the use of the midpoint distribution in the JSD results in a symmetric and normalized measure. When the distributions $p$ and $q$ are essentially the same over the $K$ states, we end up with a very small value for the $JSD(p||q)$ and thus we get a value very close to one for the BS-Index. At the other extreme, when the geometries of the $K$ states and their frequencies are poorly aligned, reflecting substantial density imbalances across states, the $JSD(p||q)$ increases toward one, yielding a correspondingly small BS-Index value.
@@
It is important to note that, unlike the Boltzmann-Shannon Interaction Entropy ––– where two distinct partitions are used to define $p$ and $q$ ––– the BS-Index uses the same categorical bins determined by the given clustering or labeling. The two distributions $p$ and $q$ are instead constructed using different underlying measures: one frequency-based and one geometry-based. Further, while the frequency-based distribution is common and self-explanatory, the geometric measure-based distribution may be defined or estimated in a number of meaningful ways. However, it requires that clusters corresponds to finite subsets of $\mathcal{X}$ within a bounded domain $\mathcal{D}\subset\mathbb{R}^{d}$ for some $d\in\mathbb{N}$, ensuring that their geometric measures can be properly normalized.
@@
The most obvious measures arise in the case of geometrically separable clusters that allow a true partition of $\mathcal{D}$ into $K$ disjoint regions, $R_{1},R_{2},...,R_{K}$, such that their union equals $\mathcal{D}$. Here, $q_{i}$ may simply be the $d$-dimensional volume of the region $R_{i}$. However, it is very common to have data where no clear bounded domain is known a priori; in that case, there are many approaches that could be taken. We will not rehash the entire discussion of this challenge that was addressed in detail in previous work diggans2022geometric; diggans2023boltzmann; instead, we simply suggest one measure that works in all cases including the one described above.
@@
Since we do not want to enforce geometric separability of the clusters by label, we will rely on the powerful connection of Singular Value Decomposition (SVD) to geometry. Given labeled data, as stated above, we consider the rectangular matrices of each labeled cluster one at a time, where rows of each matrix are the $d$-dimensional vectors in a given labeled subset. We compute the SVD of each subset and multiply the $d$ singular values to get a geometric quantity that represents the volume of the set in terms of the product of the variation in each principal component direction. We chose this as our suggested geometric measure because it is defined for all partitions, even when concepts, like the convex hull of the subsets, might overlap and be skewed by outliers.
@@
The goal of the proposed index is to provide a normalized measure of how close the clustering of the data is to an equal density distribution. Coarse-grained states in real data are rarely equal in their geometric extent or complexity. The Boltzmann–Shannon Index offers an immediate, parameter-free diagnostic of the degree to which a given partition achieves balanced density across clusters, a property that is highly desirable in exploratory data analysis, yet invisible to conventional validity indices.
IV Results
@@
To illustrate the behavior and interpretability of the Boltzmann–Shannon Index, we evaluate its performance across a sequence of increasingly realistic scenarios, beginning with simple synthetic data and progressing to resource-allocation problems.
IV.1 Two-cluster example: frequency versus geometry reversal
@@
To build immediate intuition for the Boltzmann–Shannon Index, we begin with the simplest non-trivial setting: a data set partitioned into exactly two clusters. Let the empirical frequency distribution be:
$p=(\,p_{1},\;p_{2}\,)\;=\;(\alpha,\;1-\alpha),\qquad\alpha\in[0,1].$ (2)
@@
Thus, the first cluster contains proportion $\alpha$ of the points, and the second one the remaining $1-\alpha$. For the geometric distribution $q$, we deliberately choose the reversed ordering:
$q=(\,q_{1},\;q_{2}\,)\;=\;(1-\alpha,\;\alpha).$ (3)
@@
This corresponds, for example, to the situation in which the larger cluster (in population) occupies the smaller volume in feature space, and the smaller cluster is spread over a much larger volume; a classic form of severe density imbalance. Such patterns frequently arise in real-world data, for instance in anomaly-detection settings where rare events (e.g., network intrusions or equipment faults) occupy broad, diffuse regions of the state space while normal behavior is highly concentrated. The Boltzmann-Shannon Index for $p$ and $q$ is then defined as:
$\operatorname{BSI}(\alpha)$ $=$ $1-\operatorname{JSD}(p\|q)$ (4)
$=$ $1-\frac{1}{2}\Bigl[D_{\mathrm{KL}}\bigl(p\|m\bigr)+D_{\mathrm{KL}}\bigl(q\|m\bigr)\Bigr]$
@@
Thus, under this deliberate frequency-geometry reversal, the Boltzmann-Shannon Index (BSI) reduces precisely to the entropy of a Bernoulli random variable with success probability $\alpha$, or, equivalently, to the mutual information between the cluster label and an ideal uniform geometric prior.
@@
Figure 2: Two-cluster reversal example. Population frequency $p=(\alpha,1-\alpha)$ and geometric spread $q=(1-\alpha,\alpha)$. The Boltzmann–Shannon Index (solid blue curve) attains its maximum value of 1 at $\alpha=0.5$, corresponding to perfect frequency–geometry alignment, and decreases towards 0 as $\alpha\to 0^{+}$ (or $\alpha\to 1^{-}$), where the two distributions are completely inverted and nearly all points occupy the geometrically smaller (or larger) region.
@@
As showed in Fig. 2, which displays $\operatorname{BSI}(\alpha)$, when population and geometry are perfectly matched ($\alpha=0.5$), the index attains its theoretical maximum of 1; conversely, when population and geometry are perfectly inverted ($\alpha\to 0$ or $\alpha\to 1$), the index correctly approaches 0, signaling extreme density imbalance. This elementary two-cluster example already reveals the core contribution of the Boltzmann–Shannon Index: it rewards partitions in which cluster sizes are proportional to the geometric volume (or effective spread) they occupy, and penalizes the opposite situation.
@@
The remaining experiments in this section will show that the same intuitive pattern persists across synthetic mixtures, benchmark datasets and socio-economic resource-allocation problems.
IV.2 Synthetic Gaussian mixtures

### v3

# Boltzmann-Shannon Index: A Measure of Density Balance in Clustered Data

@@ Abstract
The Boltzmann–Shannon Index (BSI) for clustered continuous data is introduced as a normalized measure that captures the relationship between geometry-based and frequency-based probability distributions defined over the clusters. In essence, it quantifies the similarity across densities of the clusters, which are defined by a given labeling. This labeling may originate from a geometric partitioning of the state space itself, but need not in general. We illustrate its performance on synthetic Gaussian mixtures, the Iris benchmark data set, and a high-imbalance resource-allocation scenario, showing that the BSI provides a coherent assessment in cases where traditional metrics give incomplete or misleading signals. Moreover, in the resource-allocation setting where equal density may be associated with a “fair” distribution, we demonstrate that BSI not only detects inequality with high sensitivity, but also offers a numerically smooth measure that can be easily embedded in optimization frameworks as a regularization term for modern policy-making. Finally, the BSI also offers a new measure of the effectiveness for a given symbolic representation, i.e. coarse-grain states, for continuous-valued data recorded from complex dynamical systems.
@@ I Introduction
Entropy, as a measure of disorder, was the cornerstone of the foundation of statistical mechanics. Although in that context it was defined in terms of volumes of phase space that contain a density of microscopic configurations of a given macrostate for a thermodynamic system boltzmann1866mechanische, it makes clear that “the logarithmic connection between entropy and probability was first stated by Ludwig Boltzmann in his kinetic theory of gases” planck1914theory.
It was almost a century later that Claude Shannon re-framed the concept of entropy in the more abstract context of information theory, where he defined it as a quantitative measure of uncertainty over a discrete probability distribution used in symbolic communication. He established a framework using this concept for understanding the fundamental limits of signal processing and data compression, with a focus on the concept of channel capacity shannon1948mathematical. And, while Shannon’s formulation is similarly a cornerstone of modern data science, it was designed for applications in discrete symbolic state spaces.
All applications to continuous state spaces, therefore, involve specific assumptions about smooth probability mass functions (outlined in the final chapter of his seminal work shannon1948mathematical). However, as E. T. Jaynes emphasized in his reformulation of information theory through a more Bayesian lens jaynes1957information, these assumptions are not self-evident. Most importantly, Shannon entropy is always defined relative to an underlying measure, and so any approach based on standard Riemann sums would correspond to choosing a uniform prior over the bins, i.e. arbitrarily chosen states. Or as Jaynes put it:
@@
When analyzing data in continuous state spaces, frequency of observation is not the only measure of interest; another important factor to consider is the geometry of the distribution of data. Data points occupy space, form shapes, and concentrate in regions of varying density. For this reason, recent work explored a specific form of metric entropy sinai1959notion; kolmogorov1985new, termed geometric partition entropy diggans2022geometric; diggans2025generalizing, which defines entropy directly from the spatial organization of data rather than from symbolic frequencies. Instead of using traditional histogram binning approaches, Geometric Partition Entropy (GPE) partitions the state space into coarse-grained regions so that each bin contains approximately the same number of sample points (e.g., using quantiles in one dimension). The entropy is then estimated from the distribution of proportions each state takes up within a bounded state space, providing a geometry-driven measure of uncertainty. This is akin to a Lebesgue integral formulation where the range of the CDF, $[0,1]$, is binned, as opposed to the domain in the Riemann integral of Shannon (see Fig. 1, reproduced from diggans2023boltzmann with the permission of the authors, for an illustration of the relationship between a GPE estimate and the corresponding Riemann estimate of Shannon entropy).
Figure 1: Comparison between frequency-based and geometry-based quantifications of uncertainty in one-dimensional data. An example of (a) an underlying probability distribution $\rho(x)$ together with (b) a corresponding cumulative distribution function $CDF(x)$. A traditional Riemann sum (frequency-based) estimate for entropy uses (c) a distribution defined by the differences in the empirical CDF at fixed equal length bin boundaries; the probability $p_{i}$ of the $i$-th bin is given by the height of the corresponding blue step. In contrast, geometric partition entropy (GPE) constructs a measure-based representation as in (d) by partitioning the co-domain $[0,1]$ into equal masses and mapping the boundaries back through the inverse CDF; the probability $q_{i}$ is obtained from the geometric length $l_{i}$ taken as a proportion of the total state-space length. This figure was reproduced with the permission from the authors of diggans2023boltzmann.
@@
Thus, for many complex data distributions there is clear utility in both viewpoints. In fact, many advances in information theory, e.g., the popular KNN method for estimating mutual information, have originated from the inclusion of geometry and frequency. This trend perhaps started with the work of Kozachenko and Leonenko kozachenko1987sample, and the Boltzmann-Shannon Interaction Entropy (BSIE) diggans2023boltzmann sought to highlight that earlier work. By combining the Boltzmann-inspired geometric view of GPE with Shannon’s frequentist formulation, BSIE provided a normalized measure that captured how the underlying geometry of a set of observations of continuous variables interacts with a corresponding estimated probability mass function. Like that earlier work, rather than treating shape and frequency as separate aspects of a dataset or averaging away its effects (as in KNN), BSIE sought to quantify the joint structure of a pair of related distributions (e.g., see ).
@@
While BSIE provided a measure that incorporated both the geometry and frequency of data in continuous variables, it did not directly account for scenarios in which the data are already partitioned (i.e., clustered), which is a common occurrence. In this case, labels are assigned that group the data into a number of disjoint subsets. The challenge is then not necessarily to quantify the entropy of a distribution, but rather to understand the impact a given choice of partition has on the resulting symbolic representation through the underlying geometry and density of the states.
@@
For example, it is common practice to label a dataset as imbalanced solely on the basis of the target distribution, typically when one class accounts for the vast majority of observations. But, when data lie in continuous spaces, imbalance may not be determined by frequency-based probability alone. For example, consider a scenario where the outcome variable follows a $90:10$ distribution, but the underlying state space is also partitioned in a way that reflects the same $90:10$ proportion. In such a case, the dataset should not be considered imbalanced. The frequency distribution is fully consistent with the geometric structure of the data, and the representation of data proportionally expresses the regions of interest in the feature space. True balance or imbalance therefore arises from the interaction between class frequencies and the geometry of the feature space, not from frequencies in isolation.
Here, we introduce the Boltzmann-Shannon Index (BSI), which gives a sense of how well the state space is partitioned into clusters where equal density clusters would result in unity.
@@
Unlike existing clustering validity indices, the proposed BSI captures both the geometric spread of the data and the frequency with which regions of the state space are occupied. Distance-based metrics, such as the Silhouette Score rousseeuw1987silhouettes, Davies–Bouldin Index davies2009cluster, and Calinski–Harabasz Index calinski1974dendrite, quantify compactness or separation but rely solely on pairwise distances or variance ratios, ignoring how density is distributed across clusters. Moreover, they are not normalized and so can be difficult to compare across datasets or scales. Information-theoretic measures, such as normalized mutual information strehl2002cluster, overcome some of these issues but require ground-truth labels and therefore cannot be used in fully unsupervised settings. In contrast, the Boltzmann–Shannon Index is agnostic, normalized, and uniquely incorporates both geometry-based and frequency-based probabilities, enabling it to detect imbalances in how the state space is partitioned that traditional metrics cannot reveal.
@@
We define the Boltzmann-Shannon Index in Sec. II, and apply this measure to several data sets with increasing relevance and complexity in Sec.  III. We then conclude with a summary of the work and plans for adaptation to mixed continuous/discrete datasets and offer several directions for future applications.
II Methodology
@@
For the purposes of this work, the specific algorithm used for clustering is irrelevant as the BSI operates "post-hoc" on any set of cluster labels. Its role is to quantify how effectively a given clustering or partitioning scheme organizes a continuous state space into a meaningful coarse-grained representation. Since our interest lies in evaluating a given partition rather than in how it is produced, we need only rely on the common $K$-Means clustering lloyd1982least, which partitions data into $K$ clusters by minimizing within-cluster variance.
@@
Given a data set $\mathcal{X}=\left\{x_{0},x_{1},...,x_{N}\right\}\subset\mathbb{R}^{d}$ and labels $\mathcal{L}=\left\{\ell_{0},\ell_{1},...,\ell_{N}\right\}$ where $\ell_{i}\in\left\{1,...,K\right\}$, we define the Boltzmann-Shannon Index to be one minus the Jensen Shannon Divergence (JSD) between a geometric measure-based distribution over the $K$ states of the partition, i.e. clusters, and the frequency-based distribution of the same $K$ states. If we let q represent a normalized geometric distribution and p represent the commonly used normalized frequency-based distribution (i.e. a histogram), then we define the BSI as:
$BSI(\mathcal{X};\mathcal{L})$ $=1-JSD(\textbf{p}||\textbf{q})$ (1)
$=1-\frac{D_{KL}(\textbf{p}||\textbf{m})+D_{KL}(\textbf{q}||\textbf{m})}{2},$
@@
where $D_{KL}(x||y)$ is the Kullback-Leibler divergence between distributions $x$ and $y$, and m is taken to be the midpoint distribution defined as the average of p and q, i.e. $\textbf{m}=(\textbf{p}+\textbf{q})/2$.
@@
Although the Kullback-Leibler divergence is asymmetric, the use of the midpoint distribution in the JSD results in a symmetric and normalized measure. When the distributions p and q are essentially the same over the $K$ states, we end up with a very small value for the $JSD(\textbf{p}||\textbf{q})$ and thus we get a value very close to one for the BSI. At the other extreme, when the geometries of the $K$ states and their frequencies are poorly aligned, reflecting substantial density imbalances across states, the $JSD(\textbf{p}||\textbf{q})$ increases toward one, yielding a correspondingly small BSI value.
@@
It is important to note that, unlike the Boltzmann-Shannon Interaction Entropy ––– where two distinct partitions are used to define p and q ––– the BSI uses the same categorical bins that are determined by the given clustering or labeling. The two distributions p and q are instead constructed using different underlying measures, one frequency-based and one geometry-based. Further, while the frequency-based distribution is common and self-explanatory, the geometric measure-based distribution may be defined or estimated in a number of meaningful ways. However, it requires that clusters corresponds to finite subsets of $\mathcal{X}$ within a bounded domain $\mathcal{D}\subset\mathbb{R}^{d}$ for some $d\in\mathbb{N}$, ensuring that their geometric measures can be properly normalized.
@@
The most obvious measures arise in the case of geometrically separable clusters that allow a true partition of $\mathcal{D}$ into $K$ disjoint regions, $R_{1},R_{2},...,R_{K}$, such that their union equals $\mathcal{D}$, e.g. based on a Voronoi diagram. Here, $q_{i}$ may simply be the $d$-dimensional volume of the region $R_{i}$. However, it is very common to have data where no clear bounded domain is known a priori; in that case, there are many approaches that could be taken. This challenge was addressed in detail in previous work diggans2022geometric; diggans2023boltzmann and so we will not revisit it here; instead, we simply suggest a measure that works in all cases including the partition described above.
@@
Since we do not want to enforce geometric separability of the clusters by label, we will rely on the powerful connection of Singular Value Decomposition (SVD) to geometry. Given labeled data, as stated above, we consider the rectangular matrices of each labeled cluster one at a time, where rows of each matrix are the $d$-dimensional vectors in a given labeled subset. We compute the SVD of each subset and multiply the $d$ singular values to get a geometric quantity that is the product of the variation in each of the principal component direction. We chose this as our suggested geometric measure because it is defined for all partitions, even when concepts like the convex hull of the subsets might overlap and be skewed by outliers.
@@
Coarse-grained states representing real dynamical time series data are rarely equal in their geometric extent or frequency of occurrence. One goal of the proposed index is to provide a normalized measure of how close such a clustering of data is to an equal density distribution, which often results in more expressive transition matrices. The Boltzmann–Shannon Index then offers an immediate, parameter-free diagnostic of the degree to which a given partition achieves balanced density across clusters, a property that is highly desirable in exploratory data analysis, yet invisible to conventional validity indices.
III Results
@@
To illustrate the behavior and interpretability of the Boltzmann–Shannon Index, we evaluate its performance across a sequence of increasingly realistic scenarios, beginning with simple synthetic data and progressing toward more application appropriate resource-allocation problems.
III.1 Two-cluster example: frequency versus geometry reversal
@@
To provide intuition for the Boltzmann–Shannon Index, we begin with the simplest non-trivial setting: a data set partitioned into exactly two clusters. Let the empirical frequency distribution be:
$\textbf{p}=(\,p_{1},\;p_{2}\,)\;=\;(\alpha,\;1-\alpha),\qquad\alpha\in[0,1].$ (2)
@@
Here, the first cluster contains the proportion $\alpha$ of the points, and the second one the remaining $1-\alpha$. For the geometric distribution q, we deliberately choose the reversed ordering:
$\textbf{q}=(\,q_{1},\;q_{2}\,)\;=\;(1-\alpha,\;\alpha).$ (3)
@@
This corresponds, for example, to the situation in which the larger cluster (in population) occupies the smaller volume in feature space, and the smaller cluster is spread over a much larger volume; a classic form of severe density imbalance. Such patterns frequently arise in real-world data, for instance in anomaly-detection settings where rare events (e.g., network intrusions or equipment faults) occupy broad, diffuse regions of the state space while normal behavior is highly concentrated. The Boltzmann-Shannon Index for p and q is then defined analytically as:
$\operatorname{BSI}(\alpha)$ $=$ $1-\operatorname{JSD}(\textbf{p}\|\textbf{q})$ (4)
$=$ $1-\frac{1}{2}\Bigl[D_{\mathrm{KL}}\bigl(\textbf{p}\|\textbf{m}\bigr)+D_{\mathrm{KL}}\bigl(\textbf{q}\|\textbf{m}\bigr)\Bigr]$
@@
Thus, under this deliberate frequency-geometry reversal, the Boltzmann-Shannon Index reduces precisely to the entropy of a Bernoulli random variable with success probability $\alpha$, or, equivalently, to the mutual information between the cluster label and an ideal uniform geometric prior.
@@
Figure 2: Two-cluster reversal example. Population frequency $\textbf{p}=(\alpha,1-\alpha)$ and geometric spread $\textbf{q}=(1-\alpha,\alpha)$. The Boltzmann–Shannon Index (solid blue curve) attains its maximum value of 1 at $\alpha=0.5$, corresponding to perfect frequency–geometry alignment, and decreases towards 0 as $\alpha\to 0^{+}$ (or $\alpha\to 1^{-}$), where the two distributions are completely inverted and nearly all points occupy the geometrically smaller (or larger) region.
@@
As showed in Fig. 2, $\operatorname{BSI}(\alpha)$ attains its theoretical maximum of one when population and geometry are perfectly matched ($\alpha=0.5$); conversely, when population and geometry are perfectly inverted (in either case $\alpha\to 0$ or $\alpha\to 1$), the index correctly approaches 0, signaling extreme density imbalance. This elementary two-cluster example already reveals the core contribution of the Boltzmann–Shannon Index in that it rewards partitions in which cluster sizes are proportional to the geometric volume (or effective spread) they occupy, and penalizes the opposite situation.
@@
The remaining experiments in this section will show that the same intuitive pattern persists across synthetic mixtures, benchmark datasets and socio-economic resource-allocation problems, but additional interesting details emerge.
III.2 Synthetic Gaussian mixtures